# Precision Mars Entry Navigation with Atmospheric Density Adaptation via Neural Networks

### v1

@@ 6 Conclusions
This work introduces a new approach to onboard real-time filtering for Martian entry using a neural network to estimate atmospheric density and employing a consider analysis to account for the uncertainty in this parameter. The network is trained on a simple exponential atmospheric density model, and its parameters are dynamically adapted in real-time to address any disparities between the true and estimated densities. The adaptation of the network is formulated as a maximum likelihood problem, leveraging the measurement innovations of the filter in a nonlinear fashion to identify optimal network parameters. The incorporation of a neural network enables the use of stochastic optimizers from the machine learning domain (such as Adam) within the context of the maximum likelihood adaptive estimation. Adam is highly optimized for neural networks and known for its accuracy and efficiency.
Three different filtering strategies were implemented to evaluate the effectiveness of the new adaptation scheme. An unscented Kalman filter with covariance matching, an unscented Kalman filter with density adaptation by state augmentation, and the newly presented approach were tested through a Monte Carlo analysis with different possible atmospheric profiles sampled from Mars-GRAM. The new approach outperformed the other two filters in terms of estimation error, yielding mostly unbiased and consistent estimates. Additionally, the new approach successfully adapted the onboard density model to match the atmospheric profiles provided by Mars-GRAM within an error of 0.7%.
Future work includes the incorporation of this new adaptation technique into a closed-loop simulation to evaluate the robustness of the solution to control inputs. Additionally, the modeled dynamics can be expanded to encompass the attitude of the vehicle as part of the state to leverage measurements from the gyroscope in the inertial measurement unit, potentially improving filtering performance and nonlinear adaptation.

### v2

@@ 7 Conclusions
This work introduces a new approach to onboard adaptive filtering for Martian entry using a neural network to estimate atmospheric density, and employing a consider analysis to account for the uncertainty in this parameter. The network is trained offline on a simple exponential atmospheric density model, and its parameters are dynamically adapted in real time to address any disparities between the true and estimated atmospheric densities. The adaptation of the network is formulated as a maximum likelihood problem by leveraging the measurement innovations of the filter to identify the optimal network parameters. The incorporation of a neural network in the current framework enables the use of stochastic optimizers from the machine learning domain, such as Adam, within the context of maximum likelihood adaptive estimation. These stochastic optimizers have been designed for neural networks and are known for their accuracy and efficiency.
Three filtering strategies were implemented and compared to evaluate the effectiveness of the proposed adaptation scheme. An unscented Kalman filter with covariance matching, an unscented Kalman filter with density adaptation by state augmentation, and the newly presented approach were tested through Monte Carlo analysis with different possible atmospheric profiles sampled from Mars-GRAM. The new approach outperformed the other two filters in terms of the estimation error, yielding mostly unbiased and consistent estimates. In addition, the new approach successfully adapted the onboard density model to match the atmospheric profiles provided by Mars-GRAM within an error of 0.7%.